### v1

# A Fully Analog Implementation of Model Predictive Control with Application to Buck Converters

Abstract
This paper proposes a novel approach to design analog electronic circuits that implement Model Predictive Control (MPC) policies for plants described by affine models. The combination of state-of-the-art approaches to define reduced-complexity Explicit MPC (EMPC) is employed to realize an analog circuit characterized by a limited amount of low-latency and commercially available components. The practical feasibility and effectiveness of the proposed approach are demonstrated through its application in the design of an advanced controller for DC-DC Buck converters. We formally analyze the stability of the obtained system and conduct extensive numerical simulations to demonstrate that it is capable of achieving outstanding load disturbance rejection performance, outclassing standard approaches.
1 Introduction
Model Predictive Control (MPC) is a widely employed control strategy in which, at each sampling period, the control input minimizes a given cost function, subject to input and state constraints. The MPC control framework is theoretically rich [1, 2] and particularly flexible, thanks to its inherent ability to handle constraints [3].
Despite these desirable features, the practical implementation of MPC is limited. First, the typical implementation of MPC imposes a significant overhead in terms of costs and circuit area due to the need to employ expensive digital hardware, including analog-to-digital converters (ADC), digital-to-analog converters (DAC), and a computing device, making the overall system too expensive for mass production, where a limited-budget design is often a primary requirement. Second, despite the rapid advancement in the development of novel algorithms for convex optimization [4, 5], the need to solve a quadratic programming (QP) optimization problem at each iteration renders the solution infeasible for systems exhibiting fast dynamics, when fast sampling is needed.
A popular approach, enabling considerable computational savings, consists in the computation of an explicit MPC (EMPC) control law. As shown in [6], the MPC control input is a piecewise affine (PWA) function, which can be computed offline, eliminating the need for online optimization. Since the pioneering work [7], several algorithms have been proposed to efficiently compute the EMPC control law, e.g., [8, 9, 10]. Still, the EMPC approach remains critical in budget-constrained and fast-sampled applications. On the one hand, the need for digital hardware persists; on the other hand, in many practical scenarios, the PWA function is defined by a large number of regions, whose real-time evaluation may remain critical. These issues are reported in, e.g., the work [11], which applies EMPC to Buck converters.
Alternatively, fully analog circuit approaches may be considered. A notable result in this direction is [12], which constructs an analog circuit that solves linear and quadratic optimization problems, thus directly applicable to MPC implementation. However, the applications of this method are limited due to the resulting circuit complexity.
The first contribution of this paper is a novel, completely general methodology to design a fully analog electronic circuit implementing MPC controllers for linear systems. We begin with an EMPC design and apply state-of-the-art approaches to obtain a PWA description of the controller characterized by a limited amount of polytopic regions. Next, we define the circuital implementation using commercially available low-latency components: comparators, resistors, operational amplifiers (OP-AMPs), and one multiplexer (MUX).
The second contribution is the application of the proposed approach to the control of DC-DC Buck converters. Buck converters are devices used to regulate a high, possibly fluctuating, supply voltage to a lower and constant value. The application of Buck converters ranges from biomedical applications [13] to renewable energy generation [14, 15], across electric mobility [16] and consumer electronics. The accurate regulation of the output voltage in Buck converters is fundamental for the proper operation of the devices supplied by the converter, which otherwise may be damaged or not work properly. However, the tracking accuracy of the converter is hindered by unpredictable and sudden variations of the input supply voltage (line disturbances) and of the current drawn by the load (load disturbances) [17, 18]. To improve tracking performance, several approaches have been considered. As to line disturbance attenuation, we mention the popular feedforwarding technique [19]. In contrast, the problem of attenuating load disturbance effects is more challenging. It has been addressed by methods based on the design of circuitry to modify either the feedback signal or the pulse width of the Pulse Width Modulated (PWM) signal in response to load variations; see [20, 21, 22] for recent works in this direction. However, these solutions lack solid theoretical foundations. Other approaches aim at estimating the output current disturbance to apply a proper feedforward action. Among these, we mention the extended state observer [23], generalized proportional-integral observer [24], and the load estimator-compensator [18] approaches. These solutions are often combined with optimal controller design strategies: [23] uses linear quadratic gaussian control in conjunction with the extended state observer, and [18] uses $\mathcal{H}_{\infty}$ optimal control in conjunction with the load-estimator-compensator.
[…]
To apply the proposed approach to DC-DC Buck converters, we first introduce a discrete-time mathematical model of the device and its linearization. Next, to ensure proper load disturbance rejection performance, we resort to the load estimator proposed in [18], which allows for a cheap analog implementation, and apply the proposed analog EMPC design. Thanks to the adopted regions reduction strategies, we observe a significant improvement over previous works considering EMPC design on Buck (see [11]), while ensuring a cheap circuit design. Furthermore, the proposed approach leads to a less expensive and faster implementation compared to the approach proposed in [12]. Furthermore, we provide a formal proof of global and exponential stability of the unique equilibrium point of the feedback control system, based on contraction theory.
Finally, we present a thorough simulation study. We perform extensive and accurate simulation experiments to demonstrate the feasibility and effectiveness of the proposed approach. The proposed analog implementation enables high-frequency sampling at $500\text{\,}\mathrm{kHz}$, which is compatible with modern Buck design and considerably improves previous solutions, e.g.,  [26, 27, 28, 29]. We analyze the performance of the system in the presence of parametric uncertainty in the value of the Buck components through Monte Carlo simulations, considering both ceramic and electrolytic capacitor cases. Next, we perform an accurate low-level simulation utilizing the manufacturer’s models of commercially available components, demonstrating the feasibility of the proposed method even in the presence of components non-idealities, such as the finite gain-bandwidth product of the OP-AMPs and delays of comparators and MUX. Our results indicate that the controlled system enjoys outstanding load disturbance rejection performances, outperforming standard methods.
1.1 Outline
Sec. 2 reviews the standard MPC problem formulation, fixing the notation for the rest of the paper. Sec. 3 expands on the proposed design method by proposing the efficient EMPC formulation and its analog circuit implementation. In Sec. 4 we apply the proposed approach to the relevant DC-DC Buck converter control problem: we introduce the system model, detail how to estimate unmeasurable quantities, and provide a rigorous stability guarantee. Sec. 5 depicts the obtained numerical results. Finally, Sec. 6 draws conclusions.
1.2 Notation
In the following, $x=[x_{i}]_{i=1}^{N}\in\mathbb{R}^{nN}$ is the vertical concatenation of the vectors $x_{i}\in\mathbb{R}^{n}$; $I_{n}\in\mathbb{R}^{n\times n}$ is the identity matrix; $\bm{0}_{n},\bm{1}_{n}\in\mathbb{R}^{n}$ are the all-0 and all-$1$ vectors, respectively; given $A\in\mathbb{R}^{n\times n}$, $\left\|x\right\|_{A}\doteq\sqrt{x^{\top}Ax}$ is the weighted $\ell_{2}$-norm of $x\in\mathbb{R}^{n}$ and $\left\|G\right\|_{A}$ is the corresponding induced matrix norm; $\otimes$ is the Kronecker product; $e_{n}^{(m)}\in\mathbb{R}^{n}$ is the $m$-th vector of the standard Euclidean basis of $\mathbb{R}^{n}$; $E_{n}^{(m)}\in\mathbb{R}^{n\times n}$ is the matrix with $1$s on the $m$-th subdiagonal and 0s elsewhere; $\odot$, $\oplus$, and $\bar{\cdot}$ are the AND, OR, and NOT operations on Boolean (logic) variables, respectively.
2 MPC Problem Formulation
This section reviews the MPC formulation that we will consider throughout the remainder of the paper.
Let us consider a DT affine system $\mathcal{A}$ described by
$x_{k+1}$ $=Ax_{k}+Bu_{k}+B_{\nu}\nu_{k}+b,$ (1a)
$y_{k}$ $=Cx_{k}+Du_{k}+D_{\nu}\nu_{k}+d,$ (1b)
where $x_{k}\in\mathbb{R}^{n}$ is the state vector, $u_{k}\in\mathbb{R}^{n_{u}}$ is the input, $\nu_{k}\in\mathbb{R}^{n_{\nu}}$ is a vector of exogenous disturbances, $y_{k}\in\mathbb{R}^{n_{y}}$ is the output signal under control, $A\in\mathbb{R}^{n\times n}$, $B\in\mathbb{R}^{n\times n_{u}}$, $B_{\nu}\in\mathbb{R}^{n\times n_{\nu}}$, $C\in\mathbb{R}^{n_{y}\times n}$, $D\in\mathbb{R}^{n_{y}\times n_{u}}$ and $D_{\nu}\in\mathbb{R}^{n_{y}\times n_{u}}$ are matrices describing the system dynamics, while $b\in\mathbb{R}^{n}$ and $d\in\mathbb{R}^{n_{y}}$ are constant affine terms.
We define, for each $k\geq 0$, the MPC optimal control problem for the plant (1) as
[…]
s.t. $\hat{x}_{0}=x_{k},\quad\hat{x}_{i+1}=A\hat{x}_{i}+B\hat{u}_{i}+B_{\nu}\nu_{k}+b,$ (2b)
$\hat{y}_{i}=C\hat{x}_{i}+D\hat{u}_{i}+D_{\nu}\nu_{k}+d,$ (2c)
$H_{x}\hat{x}_{i}\leq h_{x},\quad H_{u}\hat{u}_{i}\leq h_{u},$ (2d)
$i=0,\ldots,N_{p}-1.$
Here, the decision variables are the predicted state trajectory $\hat{x}=[\hat{x}_{i}]_{i=1}^{N_{p}}\in\mathbb{R}^{nN_{p}}$, the predicted output trajectory $\hat{y}=[\hat{y}_{i}]_{i=0}^{N_{p}-1}\in\mathbb{R}^{n_{y}N_{p}}$, and the input trajectory $\hat{u}=[\hat{u}_{i}]_{i=0}^{N_{p}-1}\in\mathbb{R}^{n_{u}N_{p}}$, $N_{p}$ is the prediction horizon, $J:\mathbb{R}^{n_{y}}\times\mathbb{R}^{n_{u}}\rightarrow\mathbb{R}$ is the cost function, $(H_{x},h_{u})$ and $(H_{u},h_{u})$ represent linear inequality constraints on the state and input, respectively.
We define the MPC cost function in Eq. (2a) as
[…]
$+\sum_{i=1}^{N_{p}-1}\|\hat{u}_{i}-\hat{u}_{i-1}\|_{R_{\Delta}}^{2},$ (3)
where $(u_{r},y_{r})\in(\mathbb{R}^{n_{u}},\mathbb{R}^{n_{y}})$ are constant input and output references we aim to track, and $Q,R_{\Delta}\succeq 0$, $R\succ 0$ are symmetric weighting matrices of suitable dimensions.
Assumption 1.
The triple $(u_{r},x_{r},y_{r})$ is an equilibrium of the undisturbed system (1), i.e., with $\nu_{k}=\bm{0}_{n_{\nu}}$ it holds
$x_{r}$ $=Ax_{r}+Bu_{r}+b,$ (4a)
$y_{r}$ $=Cx_{r}+Du_{r}+d.$ (4b)
The cost function (2) it is composed of three terms:
  • $\bullet$
    $\|\hat{y}_{i}-y_{r}\|_{Q}^{2}$ and $\|\hat{u}_{i}-u_{r}\|_{R}^{2}$ serve as tracking terms for the predicted output $\hat{y}_{i}$ and input $\hat{u}_{i}$ towards the reference $y_{r}$ and $u_{r}$, respectively;
  • $\bullet$
    $\|\hat{u}_{i}-\hat{u}_{i-1}\|_{R_{\Delta}}^{2}$ penalizes the variation in time of the inputs to obtain smoother predicted trajectories (see, e.g., [30, 31]).
Equality constraints (2b) and (2c) serve as MPC prediction model, which corresponds to model (1). The linear inequality constraints (2d) define lower and upper bounds on the input variables to explicitly handle saturation nonlinearities.
By assuming that all states and disturbances are available for measurement, we define the control law by using the one-step receding horizon policy, i.e., at each time instant $k$, the first optimal input $u_{k}^{*}\doteq\hat{u}_{0}^{*}$ given by the MPC problem (2) is applied to the plant (1) over the time interval $I_{k}$. This strategy implicitly defines a static state-feedback control policy $\pi:\mathbb{R}^{n_{p}}\to\mathbb{R}^{n_{u}}$, i.e.,
$u_{k}^{*}=\pi(p_{k}),$ (5)
where
$p_{k}\doteq[x_{k}^{\top},\;\nu_{k}^{\top}]^{\top}\in\mathbb{R}^{n_{p}}$ (6)
is a vector of variables acting as parameters in Eq. (2) and $n_{p}=n+n_{\nu}$.
The MPC problem (2) can be rewritten in a simpler and more compact QP form, comprising only $\hat{u}$ as decision variables. We achieve this by eliminating the constraints (2b), (2c) and rewriting Eq. (2a) as a function of $\hat{u}$ and $p_{k}$ only. First, to eliminate the variables $\hat{x}$ and $\hat{y}$ from problem (2), by recursively applying Eq. (2b), we have
$\hat{x}_{i}=A^{i}x_{k}+\sum_{j=0}^{i-1}A^{i-j-1}(B\hat{u}_{i}+B_{\nu}\nu_{k}+b).$ (7)
Then, $\hat{x}$ is expressed as a function of $\hat{u}$ and $p_{k}$ by
$\hat{x}=\begin{bmatrix}\Phi&\Gamma_{\nu}\end{bmatrix}p_{k}+\Gamma\hat{u}+\gamma,$ (8)
where
$\Phi$ $=\begin{bmatrix}I_{n}\\ \sum_{i=1}^{N_{p}}e_{N_{p}}^{(i)}\otimes A^{i}\end{bmatrix}$, (9a)
$\Gamma$ $=\begin{bmatrix}\bm{0}_{n\times N_{p}n_{u}}\\ I_{N_{p}}\otimes B+\sum_{i=1}^{N_{p}-1}E_{N_{p}}^{(i)}\otimes A^{i}B\end{bmatrix}$, (9b)
$\Gamma_{\nu}$ $=\begin{bmatrix}\bm{0}_{n\times n}\\ \sum_{i=1}^{N_{p}}\left(e_{N_{p}}^{(i)}\otimes\sum_{j=0}^{i-1}A^{j}\right)\end{bmatrix}B_{\nu},$ (9c)
$\gamma$ $=\begin{bmatrix}\bm{0}_{n\times n}\\ \sum_{i=1}^{N_{p}}\left(e_{N_{p}}^{(i)}\otimes\sum_{j=0}^{i-1}A^{j}\right)\end{bmatrix}b.$ (9d)
By Eq. (2c), we can also rewrite
$\hat{y}=\overline{C}\hat{x}+\overline{D}\hat{u}+\overline{D}_{\nu}\nu_{k}+\overline{d},$ (10)
where
$\overline{C}$ $=\begin{bmatrix}I_{N_{p}}\otimes C&\bm{0}_{n_{y}\times n}\end{bmatrix}$, $\quad\overline{D}$ $=I_{N_{p}}\otimes D,$ (11a)
$\overline{D}_{\nu}$ $=I_{N_{p}}\otimes D_{\nu},$ $\quad\overline{d}$ $=\bm{1}_{N_{p}}\otimes d.$ (11b)
Next, by defining
$\overline{Q}=I_{N_{p}}\otimes Q,$
$\overline{R}=I_{N_{p}}\otimes R,\quad\overline{R}_{\Delta}=M\otimes R_{\Delta},$
$M=\operatorname*{diag}([1,2\cdot\bm{1}_{N_{p}-2}^{\top},1]^{\top})-E_{N_{p}}^{(1)}-E_{N_{p}}^{(1)\,\top}$,
$\overline{y}_{r}=\bm{1}_{N_{p}}\otimes y_{r},$ (12)
[…]
$J(\hat{y},\hat{u})=\|\hat{y}-\overline{y}_{r}\|_{\overline{Q}}^{2}+\|\hat{u}\|_{\overline{R}+\overline{R}_{\Delta}}^{2}.$ (13)
Replacing Eqs. (8) and (10) into Eq. (13) yields
$J(\hat{y},\hat{u})=J_{u}(\hat{u})=\frac{1}{2}\hat{u}^{\top}H\hat{u}+(Fp_{k}+c)^{\top}\hat{u}$, (14)
where
$H$ $=(\overline{C}\Gamma+\overline{D})^{\top}\overline{Q}(\overline{C}\Gamma+\overline{D})+\overline{R}+\overline{R}_{\Delta},$ (15a)
$F$ $=(\overline{C}\Gamma+\overline{D})^{\top}\overline{Q}\,\overline{C}\begin{bmatrix}\Phi&\overline{C}\Gamma_{\nu}+\overline{D}_{\nu}\end{bmatrix}$, (15b)
$c$ $=(\overline{C}\Gamma+\overline{D})^{\top}\overline{Q}(\overline{C}\gamma+\overline{d}-\overline{y}_{r}).$ (15c)
Finally, we also express the linear constraints (2d) as a function of $\hat{u}$ and $p_{k}$ only. First, we compact Eq. (2d) as
$\overline{H}_{x}\hat{x}\leq h_{x},\quad\overline{H}_{u}\hat{u}\leq h_{u}$ (16)
where
$\overline{H}_{x}$ $=\begin{bmatrix}I_{N_{p}}\otimes H_{x}&\bm{0}\end{bmatrix}$, $\quad\overline{h}_{x}$ $=\bm{1}_{N_{p}}\otimes h_{x},$ (17a)
$\overline{H}_{u}$ $=I_{N_{p}}\otimes H_{u},$ $\quad\overline{h}_{u}$ $=\bm{1}_{N_{p}}\otimes h_{u}.$ (17b)
Then, replacing Eq. (8) into Eq. (16) yields the constraints
$G\hat{u}\leq w+Kp_{k},$ (18)
where
$G$ $=\begin{bmatrix}\overline{H}_{x}\Gamma\\ \overline{H}_{u}\end{bmatrix},\quad w=\begin{bmatrix}\overline{h}_{x}-\overline{H}_{x}\gamma\\ \overline{h}_{u}\end{bmatrix}$,
$K$ $=\begin{bmatrix}-\overline{H}_{x}\Phi&-\overline{H}_{x}\Gamma_{\nu}\\ \bm{0}&\bm{0}\end{bmatrix}$. (19)
Overall, the original MPC problem (2) is equivalent to the following QP (henceforth called QP-MPC):
$\min_{\hat{u}}\quad$ $\frac{1}{2}\hat{u}^{\top}H\hat{u}+(Fp_{k}+c)^{\top}\hat{u}$ (20a)
s.t. $G\hat{u}\leq w+Kp_{k}.$ (20b)
We notice that, thanks to the positive definiteness assumptions on the matrices $Q$, $R$, and $R_{\Delta}$, it holds $H=H^{\top}\succ 0$. See [32] for a detailed discussion. Therefore, for any value of the parameters $p_{k}$, the QP-MPC Problem (20) is strongly convex and admits a unique and global optimum $\hat{u}$. See, e.g., [33, 4, 34]. Consequently, the policy $\pi$ is well-defined.
Proposition 1.
Given Assumption 1, assume that $x_{r}$ and $u_{r}$ satisfy Eq. (2d), and let $p_{k}=[x_{r}^{\top},\bm{0}_{n_{\nu}}^{\top}]^{\top}$. Then, problem (20) is uniquely solved by $\hat{u}^{*}=\bm{1}_{N_{p}}\otimes u_{r}$.
Proof.
[…]
Proposition 2.
The output reference $y_{r}$ is an equilibrium of the closed-loop undisturbed system (1), (5), i.e.,
$x_{r}$ $=Ax_{r}+B\pi([x_{r}^{\top},\bm{0}_{n_{\nu}}^{\top}]^{\top})+b,$ (21a)
$y_{r}$ $=Cx_{r}+D\pi([x_{r}^{\top},\bm{0}_{n_{\nu}}^{\top}]^{\top})+d.$ (21b)
Proof.
This result is a direct consequence of Assumption 1 and Proposition 1. ∎
3 Proposed Design Approach
This section introduces the proposed MPC design. First, we discuss the application of methodologies to reduce the number of regions in the explicit formulation; next, we detail the EMPC analog circuit implementation and its component selection.
3.1 Explicit MPC Design
Let us recall the MPC static control policy introduced in Eq. (5), i.e., $u_{k}^{*}=\pi(p_{k})$. By the strong convexity of the MPC problem (20), $\pi(\cdot)$ establishes a bijection between each value of the parameters $p_{k}$ (6) and the optimal MPC input $\hat{u}_{0}^{*}\doteq u_{k}^{*}$ (by the receding horizon policy, see Sec. 2).
Assumption 2.
The values of interest for the parameters $p_{k}$ lie in a convex polytopic set $\mathcal{P}\subseteq\mathbb{R}^{n_{p}}$.
Eq. (5) is commonly denoted as Explicit MPC (EMPC) policy. Finding the analytic expression of such a policy takes the form of a multi-parametric quadratic programming (mp-QP) problem in the decision variables $\hat{u}$ and parameters $p_{k}$; see, e.g., [6]. For QP problems matching Eq. (20), the EMPC policy (5) can be computed exactly and in closed form, as stated by the following Theorem.
Theorem 1 ([6]).
Under Assumption 2, the optimal solution of the QP-MPC problem (20), at each $k\geq 0$, is given by
$u_{k}^{*}=\hat{u}_{0}^{*}=\pi(p_{k}),$ (22)
where:
  1. i)
    $\pi:\mathcal{P}\to\mathcal{U}$ is a continuous piecewise affine (PWA) function, defined over $R$ regions $\mathcal{R}_{i}$, $i=1,\ldots,R$, i.e.,
    $u_{k}^{*}=\pi(p_{k})=K_{i}p_{k}+l_{i}\quad\textrm{if}\;\;p_{k}\in\mathcal{R}_{i}.$ (23)
  2. ii)
    The regions $\{\mathcal{R}_{i}\}_{i=1}^{R}$ are full-dimensional polytopes with non-overlapping interiors, forming a partition of $\mathcal{P}$, i.e.,
    $\bigcup_{i=1}^{R}\mathcal{R}_{i}=\mathcal{P},\quad\operatorname*{int}(\mathcal{R}_{i})\cap\operatorname*{int}(\mathcal{R}_{j})=\emptyset,\;\;i\neq j.$ (24)
Replacing the MPC problem (20) with the EMPC policy (22) offers the significant advantage of being able to compute the optimal control input using a static function. As we shall see in Sec. 3.2, thanks to the PWA structure of Eqs. (23) and (24), we can realize the EMPC employing only static analog components, achieving a very fast function evaluation. Yet, its practical applicability is limited to problems of modest size. Specifically, the number of polytopic regions of $\pi(\cdot)$, as in Eqs. (23) and (24), is problem-dependent. A high number of regions may easily make EMPC not feasible both for digital and analog practical realization, since:
  • $\bullet$
    For digital/embedded implementation, a high number of regions may exceed the storage capacity of the hardware [32]. Also, identifying the region $\mathcal{R}_{i}$ containing the current parameter $p_{k}$ (to then apply the corresponding affine control law $u_{k}^{*}=K_{i}p_{k}+l_{i}$) may take more time than solving the actual MPC problem (20) using conventional algorithms for convex optimization [32].
  • $\bullet$
    For analog implementation, a high number of regions would require a large number of components, resulting in an overly bulky and expensive circuit. We present a detailed discussion on this aspect in Sec. 3.2.
The number of regions $R$ is equal to the number of possible combinations of active constraints for problem (20)  [35]. A rough upper bound on $R$ is therefore given by $R\leq 2^{q}$, where $G\in\mathbb{R}^{q\times N_{p}n_{u}}$ and $w\in\mathbb{R}^{q}$ in Eq. (20b). In most practical applications, $R$ is smaller because most of the constraint combinations are never active at optimality. For instance, when the input and state constraints in Eq. (2d) are non-coincident upper and lower bounds on given quantities, at most one of the two inequalities is active, thus halving the maximum number of regions.
In the following, our goal is to further reduce the number of regions and the number of inequalities defining them. To this aim, we perform four consecutive steps:
  • $\bullet$
    Move blocking strategy;
  • $\bullet$
    Optimal merging of adjacent regions;
  • $\bullet$
    Hyperplane separation of saturated regions;
  • $\bullet$
    Removal of trivial inequalities.
In the remainder of this section, we detail these steps.
3.1.1 Move Blocking Strategy
A widely-employed approach to reduce the complexity of the QP-MPC problem (20) is the so-called move blocking strategy [36]. Such a strategy consists in linearly mapping the input trajectory $\hat{u}$ to a smaller set of decision variables $u_{c}=[u_{c,j}]_{j=0}^{N_{c}-1}$, $N_{c}<N_{p}$. The mapping is defined as follows:
$\hat{u}=Tu_{c},\quad T\in\mathbb{R}^{N_{p}n_{u}\times N_{c}n_{u}},$ (25)
where the matrix $T$ is lower-trapezoidal and can be chosen at will to “shape” the input trajectory $\hat{u}$. A popular choice is
$T=\begin{bmatrix}I_{(N_{c}-1)n_{u}}&\bm{0}_{n_{u}\times n_{u}}\\ \bm{0}_{n_{u}\times n_{u}}&I_{n_{u}}\\ \vdots&\vdots\\ \bm{0}_{n_{u}\times n_{u}}&I_{n_{u}}\end{bmatrix}$, (26)
so that $\hat{u}$ is a piecewise constant sequence consisting of $N_{c}$ segments: the first $N_{c}-1$ ones consist of one sample only, while the last one comprises all samples from $N_{c}-1$ to $N_{p}-1$.
Eq. (25) can be directly replaced into the MPC Problem (20), obtaining its reduced version with variables $u_{c}$. Specifically, Problem (20) is recast to
$\min_{u_{c}}\quad$ $\frac{1}{2}u_{c}^{\top}\tilde{H}u_{c}+(\tilde{F}p_{k}+\tilde{c})^{\top}u_{c}$ (27a)
s.t. $\tilde{G}u_{c}\leq w+Kp_{k}.$ (27b)
where
$\tilde{H}$ $=T^{\top}HT,\quad\tilde{G}=GT$ (28a)
$\tilde{F}$ $=T^{\top}F,\quad\tilde{c}=T^{\top}c$ (28b)
We notice that all the fundamental properties of the MPC formulation (20), such as its strong convexity and the defined equilibrium, are preserved.
The move blocking strategy translates into a reduction of the number of regions $R$ in the EMPC policy (22). In fact, the number of decision variables in the MPC problem (20) has now reduced from $N_{p}n_{u}$ to $N_{c}n_{u}$. Without move blocking, at most $n_{u}N_{p}$ constraints can be active at the optimum:
$R\leq\sum_{i=0}^{n_{u}N_{p}}\binom{q}{i}$. (29)
With move blocking ($N_{c}<N_{p}$), a new upper bound on the number of regions is:
$R\leq\sum_{i=0}^{n_{u}N_{c}}\binom{q}{i}<\sum_{i=0}^{n_{u}N_{p}}\binom{q}{i}$. (30)
For more details on this aspect, we refer the reader to [37].
With this modification, the optimal solution to the QP is not preserved; however, a suitable tuning of the parameter $N_{c}$ can always be taken to ensure a negligible decrease in the closed-loop performance.
3.1.2 Optimal Merging of Adjacent Regions
In the EMPC policy (22), regions sharing the same affine control law can be merged together, provided that such a merging produces a convex polytope [32]. A necessary condition for merging is, thus, that the latter regions are adjacent, i.e., each couple of regions shares a common facet.
An optimal merging algorithm, based on branch and bound search, was proposed in [38], allowing to obtain the minimal number of regions. In the subsequent design, we rely on this technique, as implemented in the Multi-Parametric Toolbox (MPT) for Matlab [39].
3.1.3 Hyperplane Separation of Saturated Regions
The MPC formulation (2) typically accounts for lower and upper bounds on the input, i.e., (2d) includes constraints of the kind $\underline{u}\leq\hat{u}_{i}\leq\overline{u}$ for some $\underline{u},\overline{u}\in\mathbb{R}^{n_{u}}$. In this case, the EMPC policy naturally leads to several regions where the control law is constantly equal to either $\underline{u}$ or $\overline{u}$. Henceforth, such regions are referred to as saturated regions; the other regions, in contrast, are called unsaturated regions.
Denoting with $I_{\mathrm{lb}}$ and $I_{\mathrm{ub}}$ the sets of indices of the regions saturated at the lower and upper bound, respectively, we define
$\underline{\mathcal{S}}=\bigcup_{i\in I_{\mathrm{lb}}}\mathcal{R}_{i},\quad\overline{\mathcal{S}}=\bigcup_{i\in I_{\mathrm{ub}}}\mathcal{R}_{i}.$ (31)
Also, $I_{\mathrm{unsat}}=\{1,\ldots,R\}\smallsetminus(I_{\mathrm{lb}}\cup I_{\mathrm{ub}})$ is the set of indices of the unsaturated regions. Moreover, we define the union of unsaturated regions $\mathcal{R}_{\mathrm{unsat}}$ as
$\mathcal{R}_{\mathrm{unsat}}\doteq\bigcup_{i\in I_{\mathrm{unsat}}}\mathcal{R}_{i}.$ (32)
By continuity of the EMPC policy, as in Theorem 1, it holds that $\underline{\mathcal{S}}\cap\overline{\mathcal{S}}=\emptyset$. Moreover, by Theorem 1, since the polytopes $\mathcal{R}_{i}$ do not overlap, it holds that $\mathcal{P}=\underline{\mathcal{S}}\cup(\mathcal{R}_{\mathrm{unsat}})\cup\overline{\mathcal{S}}$, $\operatorname*{int}(\mathcal{R}_{\mathrm{unsat}})\cap\operatorname*{int}(\underline{\mathcal{S}})=\emptyset$, $\operatorname*{int}(\mathcal{R}_{\mathrm{unsat}})\cap\operatorname*{int}(\overline{\mathcal{S}})=\emptyset$.
It is worth noticing that the sets $\underline{\mathcal{S}}$ and $\overline{\mathcal{S}}$ are, in general, non-convex and, possibly, non-connected.
The EMPC policy can then be rewritten as follows:
[…]
To reduce the complexity of the EMPC policy $\pi$ in Eq. (33), we can think of removing the saturated regions and replace them by means of a function $\sigma:\mathcal{P}\to\mathbb{R}$ which satisfies
$\sigma(p)<0,\;\;\forall\,p\in\underline{\mathcal{S}},\quad\sigma(p)>0,\;\;\forall\,p\in\overline{\mathcal{S}}.$ (34)
The function $\sigma$ is called separation function, as it actually separates the two sets $\underline{\mathcal{S}}$ and $\overline{\mathcal{S}}$, according to its sign.
Therefore, we can equivalently rewrite Eq. (33) as
[…]
The formulation (35) involves a reduced number of regions, thanks to the removal of saturated regions. This comes at the price of introducing the separation function $\sigma$.
To make $\sigma$ require less memory/analog components to be implemented with respect to the removed saturated regions, we must ensure that $\sigma$ is sufficiently “simple”. Thus, in the following, we seek an affine separation function (i.e., a hyperplane), in the form $\sigma(p)=a^{\top}p+b$. To this aim, we present two Theorems, establishing existence conditions and how to compute such a function.
Theorem 2.
Let $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ be the unions of two sets of polytopes, such that $\mathcal{S}_{1}\cap\mathcal{S}_{2}=\emptyset$. Let $V_{1}$ and $V_{2}$ be the sets containing the vertices of the polytopes forming $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, respectively. Finally, let $\sigma(x)=a^{\top}x+b$ be an affine function.
Then, $\sigma$ separates $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, i.e.,
$\sigma(x)<0,\;\;\forall\,x\in\mathcal{S}_{1},\quad\sigma(x)>0,\;\;\forall\,x\in\mathcal{S}_{2},$ (36)
if and only if
$\sigma(v)<0,\;\;\forall\,v\in V_{1},\quad\sigma(v)>0,\;\;\forall\,v\in V_{2}.$ (37)
Proof.
The $\Rightarrow$” part is immediate, since $V_{1}\subset\mathcal{S}_{1}$ and $V_{2}\subset\mathcal{S}_{2}$. For the $\Leftarrow$” part, let us drop the subscripts $1$ and $2$, to consider both cases in Eqs. (36) and (37) at the same time. Consider $\lambda\in\mathbb{R}^{|V|}$ such that $\lambda\geq 0$ and $\sum_{i=1}^{|V|}\lambda_{i}=1$, and associate each $\lambda_{i}$ to a vertex $v_{i}\in V$, $i=1,\ldots,|V|$. Then, by Eq. (37), it holds that
$\lambda_{i}(a^{\top}v_{i}+b)\gtreqless 0,\quad\forall\,i\in\{1,\ldots,|V|\}$
$a^{\top}\sum_{i=1}^{|V|}\lambda_{i}v_{i}+b\sum_{i=1}^{|V|}\lambda_{i}\gtrless 0$
$a^{\top}x+b\gtrless 0,$ (38)
where $x=\sum_{i=1}^{|V|}\lambda_{i}v_{i}\in\operatorname*{conv}(V)$, by definition of convex hull. Since $\mathcal{S}\subseteq\operatorname*{conv}(V)$, it holds that $x\in\mathcal{S}$, yielding Eq. (36). ∎
Theorem 3.
Under the hypotheses of Theorem 2, it exists an affine function $\sigma(x)=a^{\top}x+b$ that separates $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$ if and only if it exists $\epsilon>0$ such that the following QP is feasible:
$\min_{a,b}\;\;\;$ $\|\alpha\|_{2},\quad\alpha=[a^{\top},b]^{\top},$
s.t. $a^{\top}v^{(1)}+b\leq-\epsilon,\quad\forall\,v^{(1)}\in V_{1},$ (39a)
$a^{\top}v^{(2)}+b\geq\epsilon,\quad\forall\,v^{(2)}\in V_{2}.$ (39b)
Such a function is given by $\sigma(x)=a^{*\top}x+b^{*}$, where $\alpha^{*}=[a^{*\top},b^{*}]^{\top}$ is the unique and global optimum of problem (39).
Proof.
For the $\Rightarrow$” part, by Theorem 2, if $\sigma(x)$ separates $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$, then Eq. (37) holds. This means that it exists $\epsilon>0$ such that
$\sigma(v)\leq-\epsilon<0,\;\;\forall\,v\in V_{1},\quad\sigma(v)\geq\epsilon>0,\;\;\forall\,v\in V_{2}.$ (40)
Eq. (40) matches constraints (39a) and (39b); thus, a feasible solution to problem (39) exists.
For the $\Leftarrow$” part, if a feasible solution $\alpha^{*}$ to problem (39) exists, then constraints (39a) and (39b) (and, equivalently, Eq. (37)) are satisfied for $\alpha^{*}$. In turn, by Theorem 2, Eq. (36) holds. Thus, it exists $\sigma(x)=a^{*\top}x+b^{*}$ separating $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$.
By the convexity of QP (39), the feasible solution $\alpha^{*}$ is also the unique and global optimum of problem (39[33, 4]. ∎
Remark 1.
If the QP problem (39) does not admit a feasible solution, then there exists no affine function $\sigma(x)$ that can separate $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. In such cases, it is necessary to resort to more complex separating functions, e.g., polynomial ones [40]. Nevertheless, in the numerical results presented in Sec. 5, which refer to the application considered in this paper, as introduced in Sec. 4, we could always find an affine separator for the EMPC policy (35).
Hyperplane separation of saturated regions is most effective when a significant proportion of the regions are saturated. As observed in [35], the number of saturated regions is mainly affected by the tightness of input constraints $\mathcal{U}$ and the magnitude of the input weighting matrix $R$. Tighter constraints and/or a smaller $\|R\|$ result in more regions becoming saturated, thereby enhancing the method effectiveness.
3.1.4 Removal of Trivial Inequalities
The EMPC policy (22) is defined over the polytopic domain $\mathcal{P}$, i.e., the feasible set of MPC parameters $p_{k}$ (6). The set $\mathcal{P}$ is typically specified by the user, based on prior knowledge of the admissible values that $p_{k}$ can take. Consequently, any parameter $p_{k}$ considered in practice always belongs to $\mathcal{P}$. As a result, in the EMPC policy, the inequalities associated with the facets of $\mathcal{P}$ are trivial (i.e., always satisfied) and can therefore be removed.
Overall, a comprehensive description of the MPC controller is as follows:
[…]
where denote as $r_{i},s$ the Boolean signals
$r_{i}=\begin{cases}1&\textrm{if}\;\;R_{i}p_{k}\leq q_{i},\\ 0&\textrm{otherwise},\end{cases}\quad s=\begin{cases}1&\textrm{if}\;\;a^{*\top}p_{k}+b^{*}>0,\\ 0&\textrm{otherwise},\end{cases}$ (42)
and as $s_{a}\doteq\odot_{i\in I_{\mathrm{unsat}}}\overline{r}_{i}$ a Boolean signal whose value is 0 if there is one active unsaturated region and $1$ otherwise.
3.2 Circuital Implementation
This section presents the circuit implementation of the proposed MPC controller, using analog components only and starting from its explicit formulation in Eq. (41).
We implement the MPC using: (i) one multiplexer (MUX), (ii) a set of generalized adders, (iii) a set of comparators, and (iv) a small logic gate network, according to the circuit depicted in Figure 1. A sample and hold (S/H) circuit is included after the MUX to ensure proper sampling and avoid inter-sample oscillations of the control input.
Figure 1: EMPC implementation: high-level schematic
Below, we detail every stage of the design for the four main building blocks of the EMPC circuit.
3.2.1 Multiplexer
It is the main component of the proposed EMPC design. We use an analog multiplexer to implement the by-case PWA control law in Eq. (41). Specifically, we use a multiplexer with $n_{R}+2$ inputs: one for each unsaturated region plus the two saturated input signal values $\overline{u}$ and $\underline{u}$. While $\overline{u}$ and $\underline{u}$ are trivially imposed through a constant voltage, the inputs corresponding to the control law for each unsaturated region need to be computed online, given the current measured value of states and disturbances. To this end, we introduce generalized adders.
3.2.2 Generalized Adders
Generalized adders are a standard circuit in analog electronics [41], and their design is easily automated, as detailed in this section. We use the generalized adders to generate the voltages corresponding to the command inputs $v_{j}(t)\doteq K_{j}p(t)+l_{j}$, for each $j\in I_{\mathrm{unsat}}$, evaluated at time $t=kT_{s}$. Specifically, the inputs to the generalized adders are the voltage measures of states and disturbances, $p(t)$, and a constant voltage $V_{\textrm{batt}}$ used to implement the constant offset $l_{j}$; while the $j$-th generalized adder output is the voltage $v_{j}$. Denoting $p_{n_{p}+1}(t)=V_{\textrm{batt}}$, $K_{j,i}$ the $i$-th element of vector $K_{j}$, and $K_{j,n_{p}+1}\doteq l_{j}$, we equivalently rewrite $v_{j}(t)$ as
$v_{j}(t)$ $=\sum_{i=1}^{n_{p}+1}K_{j,i}\,p_{i}(t)$ (43a)
$=\sum_{k=1}^{N_{j}}g_{j,k}^{(+)}\,p_{j,k}^{(+)}-\sum_{k=1}^{M_{j}}g_{j,k}^{(-)}\,p_{j,k}^{(-)},$ (43b)
where $N_{j},M_{j}$ ($N_{j}+M_{j}=n_{p}+1$) is the number of inputs with positive gain (i.e., $g_{j,k}^{(+)}=K_{j,i}\geq 0$) and negative gain (i.e., $g_{j,k}^{(-)}=-K_{j,i}>0$), respectively, $p_{j,k}^{(+)}=p_{i}$ for indices $i$ such that $K_{j,i}\geq 0$ and $p_{j,k}^{(-)}=p_{i}$ for indices $i$ such that $K_{j,i}<0$. For notational simplicity, we drop the adder index $j$ in the following.
Let us define the total positive and negative gains:
$K^{(+)}=\sum_{k=1}^{N}g_{k}^{(+)}\,\,\,\,\,\text{and}\,\,\,\,\,K^{(-)}=\sum_{k=1}^{M}g_{k}^{(-)}.$ (44)
We identify two topologically distinct circuits for the considered design, depending on whether $K^{(+)}\geq K^{(-)}+1$ or $K^{(+)}<K^{(-)}+1$. In the first scenario, the final circuit implementation is shown in Fig. 2, considering the presence of the resistance $R_{N+1}^{(-)}$ (in green) only, connected to the positive input of the OP-AMP $\textrm{OP}^{\textrm{add}}$. This resistor serves to increase the overall gain of the negative side. In this condition, if we ensure that
$G_{f}+\sum_{k=1}^{N+1}G_{k}^{(-)}=\sum_{j=1}^{M}G_{j}^{(+)},$ (45)
where $G_{f}=1/R_{f}$, $G_{k}^{(-)}=1/R_{k}^{(-)}$ and $G_{j}^{(+)}=1/R_{j}^{(+)}$, we obtain that
$g_{k}^{(-)}=\frac{G_{k}^{(-)}}{G_{f}}~~~~\text{and}~~~~g_{j}^{(+)}=\frac{G_{j}^{(+)}}{G_{f}}.$ (46)
Therefore, fixed a value for $R_{f}$, we can trivially obtain $R_{k}^{(-)}$ and $R_{k}^{(+)}$ from Eq. (46) and, next, $R_{N+1}^{(-)}$ is obtained using Eq. (45). Similarly, if $K^{(+)}<K^{(-)}+1$, we increase the overall gain of the positive side by adding a dummy resistance $R_{M+1}^{(+)}$ to the positive input of $\textrm{OP}^{\textrm{add}}$, as shown in Fig. 2 (blue resistor). In this case, if
$G_{f}+\sum_{k=1}^{N}G_{k}^{(-)}=\sum_{j=1}^{M+1}G_{j}^{(+)},$ (47)
we obtain, as for the previous case, the design equations in (46), from which we can compute $R_{k}^{(-)}$ and $R_{k}^{(+)}$. Finally, $R_{M+1}^{(+)}$ is computed using Eq. (47).
Figure 2: $j$-th region law: generalized adder. The blue resistor is used if $K_{j}^{(+)}<K_{j}^{(-)}+1$, while the green one is used when $K_{j}^{(+)}\geq K_{j}^{(-)}+1$
Adders constitute the most expensive part of the device, as they require OP-AMPs for the implementation. Still, the overall number of required adders coincides with the number of unsaturated regions, which, is reduced thanks to the techniques adopted in Sec. 3.1.
3.2.3 Comparators
We use a set of comparators to obtain Boolean signals $s$ and $r_{i}$, for $i\in I_{\mathrm{unsat}}$, as defined in Sec. 3.1. Let us define the Boolean signal
$r_{ij}=\begin{cases}1&\textrm{if}\;\;-[R_{i}]_{j}p_{k}+[q_{i}]_{j}\geq 0,\\ 0&\textrm{otherwise},\end{cases}$ (48)
where $R_{i}]_{j}$ and $q_{i}]_{j}$ denote the $j$-th row of matrix $R_{i}$ and the $j$-th element of vector $q_{i}$, respectively. Then, given $N_{i}$ number of inequality constraints defining region $i$, for all $i\in I_{\mathrm{unsat}}$, we have $r_{i}=r_{i,1}\odot\dots\odot r_{i,N_{i}}$ is easily realized through logic AND gates. Obtaining each of the Boolean signals $s$ and $r_{ij}$ for $i\in I_{\mathrm{unsat}}$, $j=1,\dots,N_{i}$, requires evaluating a set of inequalities of the kind
$\alpha_{1}i_{L,\mathrm{sns},k}+\alpha_{2}\hat{v}_{C,k}+\alpha_{3}\hat{i}_{\mathrm{o},k}+\alpha_{4}v_{\mathrm{in},k}+\alpha_{5}V_{0}\geq 0,$ (49)
where $V_{0}$ is an arbitrary fixed positive voltage, $\alpha_{1},\dots,\alpha_{5}$ are gains depending on $R_{i},a^{*},q_{i}$ and $b^{*}$, whose sign is unknown a-priori.
We realize the comparison altogether with the sums in Eq. (49) through the circuit topology in Fig. 3. The selected circuit compares the voltage on the “++” terminal with that on the “-” terminal, which are obtained through voltage dividers.
Figure 3: Comparator topology (case 2 positive gains and 3 negative ones)
To this end, we first rewrite (49) as
$\sum_{i\in I_{p}}\gamma_{i}z_{i}\geq\sum_{i\in I_{m}}\gamma_{i}z_{i},$ (50)
where $I_{p},I_{m}$ are the set of indices for which $\alpha_{i}>0$ and $\alpha_{i}\leq 0$, respectively, $\gamma_{i}=0.7|\alpha_{i}|/\max_{i=1,\dots,5}|\alpha_{i}|$, and $z_{i}\in\{i_{L,\mathrm{sns},k},\hat{v}_{C,k},\hat{i}_{\mathrm{o},k},v_{\mathrm{in},k},V_{0}\}$. This way, all gains for $z_{i}$ are positive and less than one, thus enabling the use of inexpensive voltage dividers for comparisons, eliminating the need for additional OP-AMPs.
Consider the networks of $|I_{p}|+1$ and $|I_{m}|+1$ resistors, respectively on the “++” and “-” terminal of the comparator. Let $R_{g}$ be the resistor connected to ground, $R_{i}^{\mathrm{cmp}}$ be the resistor associated with input $z_{i}$, and $G_{g}^{\mathrm{cmp}}=1/R_{g}^{\mathrm{cmp}},G_{i}^{\mathrm{cmp}}=1/R_{i}^{\mathrm{cmp}}$ be the associated conductances. Then the voltages on the two terminals are
$v_{+}=\sum_{i\in I_{p}}\frac{(G_{g}^{\mathrm{cmp}}+\sum_{j\neq i}G_{j}^{\mathrm{cmp}})^{-1}}{R_{i}^{\mathrm{cmp}}+(G_{g}^{\mathrm{cmp}}+\sum_{j\neq i}G_{j}^{\mathrm{cmp}})^{-1}}z_{i}$ (51)
and similarly for $v_{-}$ with the sum over $i\in I_{m}$. Imposing the gains match
$\gamma_{i}=\sum_{i\in I_{p}}\frac{(G_{g}^{\mathrm{cmp}}+\sum_{j\neq i}G_{j}^{\mathrm{cmp}})^{-1}}{R_{i}^{\mathrm{cmp}}+(G_{g}^{\mathrm{cmp}}+\sum_{j\neq i}G_{j}^{\mathrm{cmp}})^{-1}}$ (52)
is equivalent to
$\gamma_{i}(G_{g}^{\mathrm{cmp}}+\sum_{i\in I_{p}}G_{i}^{\mathrm{cmp}})=G_{i}^{\mathrm{cmp}},\quad i\in I_{p}.$ (53)
For any fixed resistor $R_{g}$, (53) is a linear system of $|I_{p}|$ equations in $|I_{p}|$ unknowns, which we recast in matrix form as:
$\begin{bmatrix}\gamma_{1}-1&\gamma_{1}&\dots&\gamma_{1}\\ \gamma_{2}&\gamma_{2}-1&\dots&\gamma_{2}\\ \vdots&\vdots&\ddots&\vdots\\ \gamma_{|I_{p}|}&\gamma_{|I_{p}|}&\dots&\gamma_{|I_{p}|}-1\end{bmatrix}\!\begin{bmatrix}G_{1}^{\mathrm{cmp}}\\ G_{2}^{\mathrm{cmp}}\\ \vdots\\ G_{|I_{p}|}^{\mathrm{cmp}}\end{bmatrix}\!$= (54)
$\quad=\!-\!G_{g}^{\mathrm{cmp}}\begin{bmatrix}\gamma_{1}\\ \gamma_{2}\\ \vdots\\ \gamma_{|I_{p}|}\end{bmatrix}\!$.
We notice that the system matrix is always full-rank, thus admits a unique solution for the conductances $G_{i}^{\mathrm{cmp}}$. Then, resistors $R_{i}^{\mathrm{cmp}}$ are easily obtained by taking the inverse. The very same procedure applies to the network at the “-” terminal.
Overall, the number of comparators is upper-bounded by $1+\sum_{i\in I_{\mathrm{unsat}}}N_{i}$, which is often effectively optimized through the removal of trivial inequalities which we detailed in Sec. 3. This number may be further reduced if different regions share the same inequality or use opposite ones (i.e., two or more polytopes defining the PWA region share some facet); in such scenarios, we may reuse comparators from other regions, eventually introducing a NOT gate.
3.2.4 Logic Network
Table 1: Truth table implementing the logic function for the region/saturation selection. It results in the function (55).
$s$ $r_{1}$ $r_{2}$ $q_{1}$ $q_{0}$
0 0 0 1 0
0 0 1 0 1
0 1 0 0 0
0 1 1 X X
1 0 0 1 1
1 0 1 0 1
1 1 0 0 0
1 1 1 X X
A network of logic gates is needed to drive the selection signal $q_{k}$ for the MUX, using the Boolean signals $r_{i}$ and $s$. These can be easily designed using standard logic function optimization methods, such as Karnaugh maps. Table 1 shows the truth table of the design associated with this simple digital circuit for the considered case study, which consisted of two unsaturated regions only, while Eq. (55) shows the resulting logic expression, directly implementable through logic gates.
$q_{0}=\overline{r_{1}}\odot(s\oplus r_{2}),~~~~q_{1}=\overline{r_{1}\oplus r_{2}}.$ (55)
3.2.5 Comparison with the literature
[…]
4 Application to DC-DC Buck Converters
In this section, we apply the proposed approach to design a fully analog MPC Buck converter.
4.1 Buck Converter Mathematical Model
Let us consider a Buck converter operating at a fixed switching frequency $f_{\mathrm{sw}}=\frac{1}{T}$. With reference to the circuit diagram of the converter depicted in Figure 4, we denote as $i_{L}(t)$ the inductor current, $v_{C}(t)$ the capacitor voltage, $d(t)$ the duty cycle of the square wave voltage $v_{\mathrm{sq}}(t)$, $i_{o}(t)$ the drawn output current, $V_{\mathrm{in}}(t)$ the supply voltage, and $v_{o}(t)$ the output voltage to be regulated. We rely on the first principles of physics to relate the above quantities and derive the mathematical model of the Buck converter.
We describe the Buck converter as the cascade of two subsystems (see Figure 4).
[…]
The ZOH samples $d(t)$ at each time instant $t=kT$, $k\in\mathbb{Z}_{\geq 0}$, and holds it constant at the value $d_{k}\doteq d(kT)$ over the $k$-th switching period $I_{k}\doteq[kT,(k+1)T)$. We denote the ZOH output signal as
$d_{\mathrm{ZOH}}(t)=d_{k},\quad\forall\,t\in I_{k}.$ (56)
Then, $d_{\mathrm{ZOH}}(t)$ is compared to a sawtooth wave, with fixed frequency $f_{\mathrm{sw}}$ and ranging in $\unitV[0,1]\;\unit${V}. This operation is performed by means of a comparator. When the sawtooth value is below $d_{\mathrm{ZOH}}(t)$, the comparator outputs a voltage $V_{H}$; conversely, when the sawtooth value is larger than $d_{\mathrm{ZOH}}(t)$, the comparator outputs a voltage $V_{L}$. The comparator output signal drives the two switches. One switch is connected to $V_{\mathrm{in}}(t)$ and is called the high side; the other is connected to the ground and called the low side. The switches are driven in anti-synchronized mode, i.e., when the comparator outputs $V_{H}$, the high side is open and the low side is closed; while if the comparator outputs $V_{L}$, the high side is closed and the low side is open. As a result, the output of subsystem $\mathcal{S}_{1}$ is a square wave $v_{\mathrm{sq}}(t)$ with duty cycle $d(t)$ and amplitude $V_{\mathrm{in}}(t)$.
Accordingly, $\mathcal{S}_{1}$ is a nonlinear block that we describe according to the following definition:
$\mathcal{S}_{1}(d(t),V_{\mathrm{in}}(t))\doteq\begin{cases}V_{\mathrm{in}}(t),&t\in[kT,(k+d_{k})T),\\ 0,&t\in[(k+d_{k})T,(k+1)T).\end{cases}$ (57)
The subsystem $\mathcal{S}_{2}$ is dynamical and linear-time-invariant (LTI). Its inputs are the square wave voltage $v_{\mathrm{sq}}(t)$ from $\mathcal{S}_{1}$ and the output current $i_{o}(t)$. The output of $\mathcal{S}_{2}$ is the controlled output voltage $v_{o}(t)$. Following the standard approach in circuit theory, we define the vector of states as the collection of inductor currents and capacitor voltages, i.e.,
$x(t)=\begin{bmatrix}i_{L}(t)\\ v_{C}(t)\end{bmatrix}$. (58)
With this choice, we obtain the following state-space model for $\mathcal{S}_{2}$:
$\mathcal{S}_{2}:\left\{\begin{aligned} \dot{x}(t)&=A_{c}x(t)+B_{c,1}i_{o}(t)+B_{c,2}v_{\mathrm{sq}}(t),\\ v_{o}(t)&=C_{c}x(t)+D_{1}i_{o}(t),\end{aligned}\right$. (59)
where
$A_{c}$ $=\begin{bmatrix}-\frac{R_{L}\parallel R_{C_{o}}}{L}&-\frac{R_{L}}{L(R_{C_{o}}+R_{L})}\\ \frac{R_{L}}{C_{o}(R_{C_{o}}+R_{L})}&-\frac{1}{C_{o}(R_{C_{o}}+R_{L})}\end{bmatrix}$, (60a)
$B_{c,1}$ $=\begin{bmatrix}\frac{R_{L}\parallel R_{C_{o}}}{L}\\ -\frac{R_{L}}{C_{o}(R_{C_{o}}+R_{L})}\end{bmatrix},\quad B_{c,2}=\begin{bmatrix}\frac{1}{L}\\ 0\end{bmatrix}$, (60b)
$C_{c}$ $=\begin{bmatrix}R_{L}\parallel R_{C_{o}}&\frac{R_{L}}{R_{L}+R_{C_{o}}}\end{bmatrix}$, (60c)
$D_{1}$ $=-R_{L}\parallel R_{C_{o}}.$ (60d)
We refer the reader to [17] for details.
Overall, the model $\mathcal{S}$ of the entire plant is a nonlinear system described by the interconnection of $\mathcal{S}_{1}$ and $\mathcal{S}_{2}$. Specifically,
$v_{o}(t)=\;$ $\mathcal{S}(d(t),i_{o}(t),V_{\mathrm{in}}(t))$ (61a)
$\doteq\;$ $\mathcal{S}_{2}(i_{o}(t),\mathcal{S}_{1}(d(t),V_{\mathrm{in}}(t))).$ (61b)
Notice that $v_{o}(t)$ depends on the $d(t)$, which we use as control input, but also on $i_{o}(t)$ and $V_{\mathrm{in}}(t)$. The latter two quantities are external signals that we can not act upon; therefore, we shall consider them as disturbances to be compensated. Moreover, we refer to $y(t)=v_{o}(t)$ as the controlled output, while $v_{o}(t)$ and $i_{L}(t)$ are the measured quantities.
According to the MPC control formulation we reviewed in Sec. 2, we need to describe the plant under control using a model of the kind $\mathcal{A}$ in Eq. (1). This section introduces suitable transformations to describe the system $\mathcal{S}$ into the desired form. Specifically, two challenges need to be addressed: first, $\mathcal{S}$ is a continuous-time model, while $\mathcal{A}$ is discrete-time; second, $\mathcal{S}$ is strongly nonlinear, while $\mathcal{A}$ is affine. We tackle these challenges through two consecutive steps: discretization and linearization.
Figure 4: Buck converter circuit diagram.
4.1.1 Model Discretization and Linearization
We select the switching period $T$ as the discretization sampling period. Henceforth, any quantity sampled at the switching time instants $t=kT$ will be denoted using the subscript $k$, e.g., $x_{k}$ will be used to denote $x(kT)$. We address discretization by integrating Eq. (59) over the $k$-th switching period $I_{k}$. Thanks to the linearity of Eq. (59), we can evaluate the model prediction $x_{k+1}$ in closed form using the following convolution integral:
$x_{k+1}=e^{A_{c}T}x_{k}\,+$
$\int_{kT}^{(k+1)T}e^{A_{c}((k+1)T-\tau)}(B_{c,1}i_{o}(t)+B_{c,2}v_{\mathrm{sq}}(t))d\tau.$ (62)
By linearity, Eq. (4.1.1) is equivalent to
$x_{k+1}=\;$ $e^{A_{c}T}x_{k}$
$+e^{A_{c}(k+1)T}\int_{kT}^{(k+1)T}e^{-A_{c}\tau}B_{c,1}i_{o}(t)d\tau$
$+e^{A_{c}(k+1)T}\int_{kT}^{(k+1)T}e^{-A_{c}\tau}B_{c,2}v_{\mathrm{sq}}(t)d\tau.$ (63)
The integrals in Eq. (4.1.1) are solved in closed form. To this aim, we introduce the following assumption:
Assumption 3.
The input voltage $V_{\mathrm{in}}(t)$ and the output current disturbance $i_{o}(t)$ remain constant over the switching periods, i.e., $V_{\mathrm{in}}(t)=V_{\mathrm{in},k}$ and $i_{o}(t)=i_{o,k}$ for all $t\in I_{k}$ and $k\geq 0$.
Remark 2.
Note that, Assumption 3 is respected with a good approximation as long as: i) the input voltage $V_{\mathrm{in}}(t)$ slowly varies (i.e., its bandwidth is much lower than $f_{\mathrm{sw}}$); ii) the output current $i_{o}(t)$ is constant almost everywhere (i.e., it is well modeled by a piecewise-constant signal, since it only changes due to sudden load connections or removals).
Under Assumption 3 and using $v_{\mathrm{sq}}(t)=\mathcal{S}_{1}(d(t),V_{\mathrm{in}}(t))$ as defined in Eq. (57), Eq. (4.1.1) simplifies to
$x_{k+1}=\;$ $e^{A_{c}T}x_{k}$
$+e^{A_{c}(k+1)T}\int_{kT}^{(k+1)T}e^{-A_{c}\tau}B_{c,1}i_{o,k}d\tau$
$+e^{A_{c}(k+1)T}\int_{kT}^{(k+d_{k})T}e^{-A_{c}\tau}B_{c,2}V_{\mathrm{in},k}d\tau.$ (64)
Solving the integrals in Eq. (4.1.1) leads to the following DT model:
$x_{k+1}=\;$ $f(x_{k},d_{k},i_{o,k},V_{\mathrm{in},k})$
$\doteq\;$ $e^{A_{c}T}x_{k}+(e^{A_{c}T}-I)A_{c}^{-1}B_{c,1}i_{o,k}$
$+e^{A_{c}T}(I-e^{-A_{c}d_{k}T})A_{c}^{-1}B_{c,2}V_{\mathrm{in},k}.$ (65)
Proposition 3.
Let $i_{o,k}=0$ and $V_{\mathrm{in},k}=\overline{V}_{\mathrm{in}}$. The system (4.1.1) admits a unique equilibrium point satisfying the output regulation task towards a desired constant reference $\overline{V}_{o}$, i.e., there exists a unique $\overline{x}\in\mathbb{R}^{n_{x}}$, $\overline{D}\in\mathbb{R}$ such that
$\overline{x}$ $=f(\overline{x},\overline{D},0,\overline{V}_{\mathrm{in}}),$ (66a)
$\overline{V}_{o}$ $=C\overline{x}.$ (66b)
Proof.
Solving Eq. (66a) for $\overline{x}$ and replacing into Eq. (66b) yields
$\overline{V}_{o}=C(I-A)^{-1}(A-A^{1-\overline{D}})A_{c}^{-1}B_{c,2}\overline{V}_{\mathrm{in}},$ (67)
with $A=e^{A_{c}T}$. Existence is guaranteed by noting that the scalar function of $\overline{D}$ on the right-hand side, defined over the interval $0\leq\overline{D}\leq 1$, admits as image the interval $0,\overline{V}_{\mathrm{in}}]$. Uniqueness is a direct consequence of the fact that such a function is strictly monotone (increasing). ∎
Remark 3.
The exact value of $\overline{D}$ can be computed, e.g., using Newton’s method to solve the nonlinear equation (67). Otherwise, taking a linear approximation of the exponential function leads to $\overline{D}=\frac{\overline{V}_{o}}{\overline{V}_{\mathrm{in}}}$, which is a good approximation in all our numerical tests.
Since the model in Eq. (4.1.1) is nonlinear, we perform linearization on it. Specifically, the DT model (4.1.1) is nonlinear in $d_{k}$ and $V_{\mathrm{in},k}$ due to the presence of the term $e^{-A_{c}d_{k}T}V_{\mathrm{in},k}$. To proceed with the required linearization, we introduce the following assumption:
Assumption 4.
The sequences $d_{k}$, $V_{\mathrm{in},k}$ are well-described by
$d_{k}=\overline{D}+\delta_{k},\quad V_{\mathrm{in},k}=\overline{V}_{\mathrm{in}}+v_{\mathrm{in},k},$ (68)
where $\delta_{k}$ and $v_{\mathrm{in},k}$ are small variations around the nominal values $\overline{D}$ and $\overline{V}_{\mathrm{in}}$, respectively.
Assumption 4 is a standard assumption in set-point regulation.Moreover, as we will demonstrate in Sec. 4.4, we are able to guarantee global exponential stability of the feedback control system when the plant is described by the non-approximated nonlinear model in Eq. (4.1.1). Therefore, Assumption 4 is only used to construct the controller design model, but is not critical for stability or performance.
Let us denote the nonlinearity of model (4.1.1) as
$g(\delta_{k},v_{\mathrm{in},k})=e^{A_{c}T}(I-e^{-A_{c}(\overline{D}+\delta_{k})T})A_{c}^{-1}B_{c,2}(\overline{V}_{\mathrm{in}}+v_{\mathrm{in},k}).$ (69)
Under Assumption 4, we can linearize Eq. (69) around the point $(\delta_{k},v_{\mathrm{in},k})=(0,0)$ as follows:
$g(\delta_{k},v_{\mathrm{in},k})\approx\;$ $g(0,0)+\frac{\partial g}{\partial\delta_{k}}(0,0)\,\delta_{k}+\frac{\partial g}{\partial v_{\mathrm{in},k}}(0,0)\,v_{\mathrm{in},k}$
$=\;$ $e^{A_{c}T}(I-e^{-A_{c}\overline{D}T})A_{c}^{-1}B_{c,2}\overline{V}_{\mathrm{in}}$
$+e^{A_{c}(1-\overline{D})T}TB_{c,2}\overline{V}_{\mathrm{in}}\,\delta_{k}$
$+e^{A_{c}T}(I-e^{-A_{c}\overline{D}T})A_{c}^{-1}B_{c,2}\,v_{\mathrm{in},k}.$ (70)
Finally, using Eq. (4.1.1) to replace $g(\delta,v_{\mathrm{in},k})$ in Eq. (4.1.1), we obtain a linearization that matches the structure of Eq. (1), where
$A$ $=e^{A_{c}T},$ (71a)
$B$ $=e^{A_{c}(1-\overline{D})T}TB_{c,2}\overline{V}_{\mathrm{in}},$ (71b)
$B_{\nu_{1}}$ $=(e^{A_{c}T}-I)A_{c}^{-1}B_{c,1},$ (71c)
$B_{\nu_{2}}$ $=e^{A_{c}T}(I-e^{-A_{c}\overline{D}T})A_{c}^{-1}B_{c,2},$ (71d)
$B_{\nu}$ $=\begin{bmatrix}B_{\nu_{1}}&B_{\nu_{2}}\end{bmatrix}$, (71e)
$b$ $=-e^{A_{c}(1-\overline{D})T}\overline{D}TB_{c,2}\overline{V}_{\mathrm{in}}.$ (71f)
Remark 4.
We highlight that, since the linearization is performed around the equilibrium point $(\overline{D},\overline{x})$, the same point is an equilibrium for the linearized system. Consequently, Assumption 1 is met.
4.2 Output Current Disturbance and State Estimation
As detailed in Sec. 2, the application of MPC design requires the measurement of the system states. Moreover, for obtaining an accurate output tracking, the disturbances should be included in the prediction. Consequently, both external disturbances and states need to be either measured or estimated. This section proposes a simple and cheap design of devices estimating $\nu(t)$ and $x(t)$.
4.2.1 Disturbances Estimation
The disturbances $\nu(t)$ are given by:
  • $\nu_{1}=i_{o,k}$. The output current drawn from the load is not directly available for measurement; therefore, it must be estimated.
  • $\nu_{2}=v_{\mathrm{in},k}$. Since $V_{\mathrm{in}}(t)$ is available for measurement, $v_{\mathrm{in},k}$ is trivially obtained as
    $v_{\mathrm{in},k}=V_{\mathrm{in},k}-\overline{V}_{\mathrm{in}}$ (72)
    and does not need to be estimated.
In the remainder of this section, we describe a method for estimating $\hat{i}_{o}(t)$ of $i_{o}(t)$. Specifically, we resort to the linear estimator proposed in [18], based on algebraic design. Specifically, given the transfer function description of the linear Buck subsystem $\mathcal{S}_{2}$ augmented with $i_{L}(t)$ as an additional output (in the Laplace domain), i.e.,
$\begin{bmatrix}v_{o}(s)\\ i_{L}(s)\end{bmatrix}=\begin{bmatrix}P_{11}(s)&P_{12}(s)\\ P_{21}(s)&P_{22}(s)\end{bmatrix}\begin{bmatrix}v_{\mathrm{sq}}(s)\\ i_{o}(s)\end{bmatrix}$, (73)
the problem of estimating the input $i_{o}$ given the measurements of $v_{o}$ and $i_{L}$ is a linear algebraic problem. In fact, it is easily seen that
$i_{o}(s)$ $=\frac{1}{P_{22}(s)}(i_{L}(s)-P_{21}(s)v_{\mathrm{sq}}(s)),$ (74a)
$v_{\mathrm{sq}}(s)$ $=\frac{1}{P_{11}(s)}(v_{o}(s)-P_{12}(s)i_{o}).$ (74b)
Replacing Eq. (74b) into Eq. (74a) and performing suitable algebraic manipulations, we obtain
$i_{o}(s)=\begin{bmatrix}-E_{1}(s)&1\end{bmatrix}\begin{bmatrix}v_{o}(s)\\ i_{L}(s)\end{bmatrix}$, (75)
where
$E_{1}(s)\doteq\frac{P_{21}(s)}{P_{11}(s)}=\frac{C(R_{L}+R_{C_{o}})s+1}{R_{L}(1+CR_{C_{o}}s)}.$ (76)
We refer the interested reader to Results 1 and 2 in Sec. III-B of [18] for a complete proof and a detailed discussion. Overall, the output current estimate $\hat{i}_{o}(t)$ is obtained by $\hat{i}_{o}(s)=\mathcal{E}(s)[v_{o}(s),i_{L,\mathrm{sns}}(s)]^{\top}$, where
$\mathcal{E}(s)\doteq\begin{bmatrix}-E_{1}(s)&g_{i_{L}}^{-1}\end{bmatrix}$. (77)
where $i_{L,\mathrm{sns}}(t)=g_{i_{L}}i_{L}(t)$ is the available measurement of the inductor current and $g_{i_{L}}\in\mathbb{R}_{+}$ is the corresponding sensor gain.
Remark 5.
The design of the estimator is entirely conducted in CT and relies on signals involved in the linear subsystem only. Consequently, the problem is solved exactly by the CT LTI filter $\mathcal{E}(s)$.
In ideal conditions, $\hat{i}_{o}(t)=i_{o}(t)$ for all $t\in\mathbb{R}_{\geq 0}$, but the presence of noise and modeling uncertainties in the plant description may hinder the quality of the estimate. Concerning noise, we notice that an appropriate printed circuit board (PCB) design easily allows us to make negligible parasitic effects perturbing the measured signals before being processed by $\mathcal{E}$. As to modeling uncertainties, we notice that only $R_{L}$, $C$, and $R_{C_{o}}$ are involved in the design, thus uncertainties on any other component are irrelevant. We study the impact of the uncertainties on $R_{L}$, $C$, and $R_{C_{o}}$ through the following sensitivity analysis:
$S^{E_{1}}_{C}(s)\doteq\frac{\partial E_{1}}{\partial C}$ $=\frac{1}{(CR_{C_{o}})^{2}}\frac{s}{(s+\frac{1}{CR_{C_{o}}})^{2}},$ (78a)
$S^{E_{1}}_{R_{C_{o}}}(s)\doteq\frac{\partial E_{1}}{\partial R_{C_{o}}}$ $=-\frac{1}{R_{C_{o}}^{2}}\frac{s^{2}}{(s+\frac{1}{CR_{C_{o}}})^{2}},$ (78b)
$S^{E_{1}}_{R_{L}}(s)\doteq\frac{\partial E_{1}}{\partial R_{L}}$ $=-\frac{1}{R_{L}^{2}}.$ (78c)
As $s\rightarrow 0$, both $|S^{E_{1}}_{C}|\rightarrow 0$ and $|S^{E_{1}}_{R_{C_{o}}}|\rightarrow 0$, indicating that uncertainty on $C$ and $R_{C_{o}}$ poorly influence the estimate $\hat{i}_{o}$ at low frequencies. Instead, $|S^{E_{1}}_{R_{L}}|$ is non-zero, indicating that the uncertainty of $R_{L}$ leads to some steady-state estimation error $|\hat{i}_{o}(t)-i_{o}(t)|$. We proceed by studying this error. Let us denote as $\hat{C},\hat{R}_{C_{o}},\hat{R}_{L}$ the nominal values of the components used to design $\mathcal{E}$, and $\hat{E}_{1}$ the transfer function $E_{1}$ as defined in Eq. (76) but evaluated at $\hat{C},\hat{R}_{C_{o}},\hat{R}_{L}$. At steady state, $v_{o}$ is constant: $v_{o}(s)=\overline{V_{o}}/s$. By the final value theorem, we obtain
$\lim_{t\rightarrow\infty}|\hat{i}_{o}(t)-i_{o}(t)|=\lim_{s\rightarrow 0}|s(\hat{i}_{o}(s)-i_{o}(s))|$
$\quad=\lim_{s\rightarrow 0}\left|s(\hat{E}_{1}(s)-E_{1}(s))\frac{\overline{V_{o}}}{s}\right|=\overline{V_{o}}\left|\frac{1}{\hat{R}_{L}}-\frac{1}{R_{L}}\right|.$ (79)
Such error is negligible when $R_{L},\hat{R}_{L}\rightarrow\infty$ or $\hat{R}_{L}\rightarrow R_{L}$. To counteract the effect of the uncertainty on $R_{L}$, we propose a robust design of $\hat{E}_{1}$ optimally selecting $\hat{R}_{L}$ to minimize the expected value of the error. Formally, let us assume that $R_{L}$ is a uniformly distributed random variable in $\underline{R}_{L},\overline{R}_{L}]$. Then our aim is to compute
$\hat{R}_{L}^{*}=\operatorname*{arg\,min}_{\hat{R}_{L}\in\mathbb{R}}\mathbb{E}\left[\left|\frac{1}{\hat{R}_{L}}-\frac{1}{R_{L}}\right|\right],\;\;R_{L}\sim U([\underline{R}_{L},\overline{R}_{L}]).$ (80)
Result 1 (Optimal estimator $\hat{R}_{L}$).
The solution to problem (80) is given by the mean resistance
$\hat{R}_{L}^{*}=\frac{1}{2}{\left(\underline{R}_{L}+\overline{R}_{L}\right)}.$ (81)
Proof.
Consider the change of coordinates $a=1/\hat{R}_{L}$ and $B=1/R_{L}$. The probability density function of $B$ is given by the change of variables formula (see Sec. 2.3 in [42]):
$f_{B}(b)=\begin{cases}\frac{1}{\overline{R}_{L}-\underline{R}_{L}}b^{-2}\;\;&\textrm{if}\;\frac{1}{\underline{R}_{L}}\leq b\leq\frac{1}{\overline{R}_{L}},\\ 0&\textrm{otherwise}.\end{cases}$ (82)
After the change of variables, the problem takes the form
$\hat{a}^{*}=\operatorname*{arg\,min}_{a\in\mathbb{R}}\mathbb{E}\left[\left|a-B\right|\right].$ (83)
It is a standard result (see, e.g., Chapter 1 in [43]) that the optimum of such a problem corresponds to the median of the distribution of $B$, i.e., $a^{*}$ satisfies
$\int_{-\infty}^{a^{*}}f_{B}(b){\rm d}b=\frac{1}{2}.$ (84)
Developing the calculations for the integral, we get
$\frac{1}{2}=\frac{1}{\overline{R}_{L}-\underline{R}_{L}}\left[-\frac{1}{b}\right]_{\overline{R}_{L}^{-1}}^{a^{*}}=\frac{\overline{R}_{L}-a^{*}}{\overline{R}_{L}-\underline{R}_{L}}.$ (85)
Solving for $a^{*}$ and taking its inverse yields the desired result. ∎
The estimator $\mathcal{E}(s)$, defined in Eq. (75), takes only the measurements of $i_{L}$ and $v_{o}$ as inputs. Moreover, $E_{1}(s)$ is a first-order filter and $i_{L}$ is not filtered. These features allow for a cheap circuit implementation of $\mathcal{E}(s)$. In the remainder of this section, we detail such an implementation, which differs from the one considered in [18].
Figure 5: Estimator circuit.
We start from the definition of $E_{1}(s)$ in Eq. (76), which can be equivalently rewritten as
$E_{1}(s)=\frac{(R_{C_{o}}\parallel R_{L})^{-1}s+\frac{1}{CR_{C_{o}}R_{L}}}{s+\frac{1}{CR_{C_{o}}}}=\frac{K_{\mathcal{E}}(s+z_{\mathcal{E}})}{s+p_{\mathcal{E}}},$ (86)
where $K_{\mathcal{E}}\doteq(R_{C_{o}}\parallel R_{L})^{-1}$, $z_{\mathcal{E}}\doteq(C(R_{C_{o}}+R_{L}))^{-1}$, and $p_{\mathcal{E}}\doteq(CR_{C_{o}})^{-1}$. Since $R_{C_{o}}+R_{L}\gg R_{C_{o}}$, we have $z_{\mathcal{E}}<p_{\mathcal{E}}$ and $E_{1}(s)$ is a high-pass filter. Consequently, we can resort to the circuital implementation of $\mathcal{E}(s)$ shown in Fig. 5. To avoid the output of the OpAmp exceeding the saturation limits, we scale the gain of $\mathcal{E}$ by a factor $g_{{i_{o}}}<1$ (0.1 in our case study), obtaining a scaled version of the current estimate $\hat{i}_{o}$. Consequently, we also multiply all gains for $\hat{i}_{o}$ by a factor $g_{i_{o}}$ in the subsequent EMPC design.
The transfer functions of the circuit in Fig. 5, from inputs $v_{o}$ and $i_{L,\mathrm{sns}}$ to output $\hat{i}_{o}$, is
$\frac{\hat{i}_{o}(s)}{v_{o}(s)}$ $=-\frac{R_{5}^{\mathcal{E}}}{R_{5}^{\mathcal{E}}}\frac{1+sR_{4}^{\mathcal{E}}C_{1}^{\mathcal{E}}}{1+sR_{5}^{\mathcal{E}}C_{2}^{\mathcal{E}}}$ (87a)
$\frac{\hat{i}_{o}(s)}{i_{L,{\rm sns}}(s)}$ $=\left(1+\frac{R_{5}^{\mathcal{E}}}{R_{5}^{\mathcal{E}}}\frac{1+s(R_{4}^{\mathcal{E}}\parallel R_{3}^{\mathcal{E}})C_{1}^{\mathcal{E}}}{1+sR_{5}^{\mathcal{E}}C_{2}^{\mathcal{E}}}\right)\cdot\frac{R_{2}^{\mathcal{E}}}{R_{1}^{\mathcal{E}}+R_{2}^{\mathcal{E}}}.$ (87b)
By imposing the conditions $g_{i_{o}}K_{\mathcal{E}}=R_{5}^{\mathcal{E}}/R_{4}^{\mathcal{E}}$, $p_{\mathcal{E}}=(R_{5}^{\mathcal{E}}C_{2}^{\mathcal{E}})^{-1}$ and $z_{\mathcal{E}}=(R_{4}^{\mathcal{E}}C_{1}^{\mathcal{E}})^{-1}$, we implement the $E_{1}(s)$ transfer function. Then, we select $R_{3}^{\mathcal{E}}$ such that $C_{1}^{\mathcal{E}}(R_{3}^{\mathcal{E}}~\parallel~R_{4}^{\mathcal{E}})=C_{2}^{\mathcal{E}}R_{5}^{\mathcal{E}}$, $R_{4}^{\mathcal{E}}=R_{2}^{\mathcal{E}}$ and $R_{5}^{\mathcal{E}}=g_{i_{L}}^{-1}g_{i_{o}}R_{1}^{\mathcal{E}}$, ensuring that $\hat{i}_{o}(s)/i_{L,\mathrm{sns}}(s)=g_{i_{L}}^{-1}g_{i_{o}}$.
4.2.2 State Estimation
All states of $\mathcal{A}$ are required as inputs to the MPC controller. Since $i_{L}(t)$ can easily be sensed (see, e.g., [18]), but the capacitor voltage $v_{C}(t)$ is not accessible due to the presence of the capacitor ESR, the estimation of $v_{C}(t)$ is needed. This section presents a simple algebraic estimator design for $v_{C}(t)$.
Starting from Kirchhoff’s current law on the output node of the Buck, we have
$i_{C}=i_{o}-i_{L}-\frac{v_{o}}{R_{L}}=\frac{v_{o}-v_{C}}{R_{C_{o}}}.$ (88)
Then, given the measurements of $i_{L}$ and $v_{o}$ and the estimate $\hat{i}_{o}$ of $i_{o}$, we define the estimator
$\hat{v}_{C}={R_{C_{o}}}(\hat{i}_{o}-i_{L})+\left(\frac{R_{C_{o}}}{R_{L}}+1\right)v_{o}.$ (89)
We highlight that, with this procedure, $\hat{v}_{C}(t)$ is obtained using only weighted sums. Consequently, we argue that compared to the conventional strategy of employing an observer, the proposed approach streamlines the design and allows for a low-cost analog implementation.
Remark 6.
In many practical applications, $R_{C_{o}}\ll R_{L}$. Under this assumption, Eq. (89) can be simplified to the straightforward relation $\hat{v}_{C}=v_{o}$, meaning that the controlled output voltage can be directly used as a reliable estimate of the state ${v}_{C}(t)$. This choice simplifies the circuit design of the system and reduces its costs both in economic terms and board area.
4.3 MPC design
We conduct MPC controller design according to the formulation in Sec. 2 and using the reduction techniques in Sec. 3.1. Specifically, within the MPC formulation in Eq. (2), we utilize the linearized Buck model described by matrices according to Eq. (71) for formulating constraints (2b) and (2c), while we include input saturation limits $\underline{u}\leq\hat{u}_{i}\leq\overline{u}$ as polytopic constraints (2d), with limits $\underline{u}=$$0\text{\,}\mathrm{V}$ and $\overline{u}=$$1\text{\,}\mathrm{V}$ representing the PWM duty-cycle limits. The selected prediction horizon is $N_{p}=5$. Regarding the weights $Q,R,R_{\Delta}$, we employ a trial-and-error procedure to optimize the qualitative observation of load and line disturbance responses. Finally, to reduce the number of regions, we apply the strategies introduced in Sec. 3.1; specifically, we select control horizon $N_{c}=2$ in the selection of the move blocking strategy map defined by the matrix in Eq. (26).
4.4 Stability Analysis
This section establishes a sufficient condition for global exponential stability of the feedback control system, given the obtained controller. We start with the following Lemma
Lemma 1.
Let $a,b\in\mathbb{R}$, $0\leq a,b\leq 1$. It holds
$(1-a)^{k}-(1-b)^{k}=\alpha_{k}(a-b)$ (90)
with $\alpha_{k}\leq 2^{k-1}$ for all $k\geq 1$.
Proof.
We prove this by induction. First, notice that the relation is trivial for $k=1$. Then, denoting $z_{a}=(1-a)^{k-1}$ and $z_{b}=(1-b)^{k-1}$, we have
$(1-a)^{k}-(1-b)^{k}=(1-a)z_{a}-(1-b)z_{b}=$ (91a)
$~=z_{a}-z_{b}+bz_{b}-az_{a}=\alpha_{k-1}(a-b)+bz_{b}-az_{a}$ (91b)
$~=a(\alpha_{k-1}-z_{a})-b(\alpha_{k-1}-z_{b})=$ (91c)
$~=a(\alpha_{k-1}-z_{b}+\alpha_{k-1}(a-b))-b(\alpha_{k-1}-z_{b})=$ (91d)
$=(a-b)(\alpha_{k-1}(a+1)-z_{b})$ (91e)
Since $\alpha_{k}=\alpha_{k-1}(a+1)-z_{b}$, $\alpha_{1}=1$ and $0\leq a,z_{b}\leq 1$, we have that $\alpha_{k}<2^{k-1}$. ∎
Result 2 (Global Exponential Stability).
Assume that $V_{\rm in},i_{o}$ are constant and possibly not assuming their nominal value. Let $P$ be the solution to the discrete Riccati equation $A^{\top}PA-P=-Q$ for some $Q\succ 0$ (which exists because $A$ is Schur stable). If
$L_{f}\doteq\left\|A\right\|_{P}+\left\|\frac{A^{2}-I}{2}A_{c}^{-1}V_{\rm in}\right\|_{P}L_{u}<1$ (92)
with
$L_{u}=\sup_{r\in I_{\rm unsat}}\left\|B_{c2}[K_{r,1},~K_{r,2}]\right\|_{P},$ (93)
the feedback control system is a contraction w.r.t. the $P$ norm, thus admits a unique global and exponentially stable (with convergence rate $L_{f}$) equilibrium point $\bar{x}$.
Proof.
Consider the closed-loop system described by
$x^{+}=F(x)=Ax+\overline{B}_{\nu}\begin{bmatrix}i_{o}\\ V_{\rm in}\end{bmatrix}+e^{A_{c}(1-u(x))T}A_{c}^{-1}B_{c,2}V_{\rm in}$ (94)
where $\overline{B}_{\nu}=[B_{\nu,1},~e^{A_{c}T}A_{c}^{-1}B_{c,2}]$ and $u(x)$ is assessed in Eq. (41). To show that (94) is a contraction with rate $L_{f}$ with respect to the $P$-norm, we have to prove that the Lipschitz constant ${\rm Lip}(F)$ of the map $F(x)$ satisfies
${\rm Lip}(F)\doteq\sup_{x_{1}\neq x_{2}}\frac{\left\|F(x_{1})-F(x_{2})\right\|_{P}}{\left\|x_{1}-x_{2}\right\|_{P}}\leq L_{f}<1.$ (95)
Let us now focus on the numerator:
$F(x_{1})-F(x_{2})=A(x_{1}-x_{2})+$ (96)
$~+\sum_{k=0}^{\infty}\frac{1}{k!}(A_{c}T)^{k}[(1-u(x_{1}))^{k}-(1-u(x_{2}))^{k}]A_{c}^{-1}B_{c,2}V_{\rm in}=$
$~=A(x_{1}-x_{2})+\sum_{k=1}^{\infty}\frac{\alpha_{k}}{k!}(A_{c}T)^{k}A_{c}^{-1}B_{c,2}V_{\rm in}(u(x_{1})-u(x_{2}))$
where $0\leq\alpha_{k}\leq 2^{k-1}$ by using Lemma 1. Therefore,
$\left\|F(x_{1})-F(x_{2})\right\|_{P}\leq\left\|A(x_{1}-x_{2})\right\|_{P}+$ (97)
$~~+\left\|\frac{A^{2}-I}{2}A_{c}^{-1}B_{c,2}V_{\rm in}(u(x_{1})-u(x_{2}))\right\|_{P}$
Finally, we have that the map ${B_{c,2}}u(x)$ is globally Lipschitz (w.r.t. the $P$-norm) thanks to the continuity and piecewise affine structure, i.e.,
$\left\|{B_{c,2}}(u(x_{1})-u(x_{2}))\right\|_{P}\leq L_{u}\left\|x_{1}-x_{2}\right\|_{P}$ (98)
with constant $L_{u}$ assessed in Eq. (93), obtained considering that for piecewise affine functions the global Lipshitz constant is the maximum of the local ones defined by the gain of each region. Notice that $L_{u}$ is independent of the coefficients $K_{r,3},K_{r,4}$ multiplying $i_{o}$ and $V_{\rm in}$ as they contribute equally to $u(x_{1})$ and $u(x_{2})$ and, thus, cancel out. This leads us to
${\rm Lip}(F)\leq\left\|A\right\|_{P}+\left\|\frac{A^{2}-I}{2}A_{c}^{-1}V_{\rm in}\right\|_{P}L_{u}$ (99)
The result follows by noting that, under the considered assumptions, the system is a contraction because ${\rm Lip}(F)<1$ and applying standard results on contraction theory [44]. ∎
Since $A$ is Schur stable, $\left\|A\right\|_{P}=\max_{i}|{\lambda_{i}(A)}|<1$ by construction. Moreover, the second term in (92) is negligible compared to the allowed margin $1-\left\|A\right\|_{P}$. In fact, despite $L_{u}$ being large in our case study because $B_{c2}=[L^{-1},0]^{\top}$ with $L$ in the order of $10^{-6}{\rm F}$, thanks to the fast sampling we have $T\approx 0$ and $A^{2}-I\approx 0$; also $A_{c}^{-1}\approx 0$ because $A_{c}$ has large eigenvalues. This simple numerical test verifies the stability of the proposed MPC controller, which can be safely adopted independently of the circuit’s initial conditions.
Remark 7.
This condition is computationally efficient: it requires computing only one discrete Riccati equation for obtaining $P$ and then $2+|I_{\rm unsat}|$ weighted norms of $n\times n$ matrices.
Remark 8.
The proposed stability result is not standard in MPC literature, and does not rely on a large prediction horizon or terminal constraints. Instead, it established a condition depending on the Lipschitz constant of the PWA controller and the plant’s structure, also exploiting the known saturation of the controller.
4.5 Comparison with the Literature
We highlight that, unlike the EMPC solution proposed in [11], the adopted strategies lead to a very reduced amount of regions, each of them defined by a few inequalities (only two regions defined by three inequalities each in our numerical case study, see Sec. 5), which makes the EMPC prone to efficient implementation.
For the proposed Buck design, compared to the analog design approach proposed in [12], we observe a reduction from 8 to 2 OP-AMPs. This comes at the cost of introducing additional logic, comparators.
5 Simulations and Results
This section presents numerical results from simulation experiments on the designed Buck feedback control system.
We consider the case study of a Buck converter defined by parameters listed in Table 2.
Table 2: Case study: Buck parameters
Parameter Value Uncertainty
[…]
We design the estimator $\mathcal{E}$ according to Sec. 4.2 and the EMPC controller according to Sec. 3 using nominal values for each component. As to the EMPC, thanks to the combination of the four proposed optimization strategies in Sec. 3, we obtain an optimal control policy $\pi(\cdot)$ defined by only $2$ unsaturated regions, and each of these regions is defined by $3$ non-trivial inequalities only (one of which is shared by the two regions). Therefore, five comparators, 2 OP-AMPs, and a four-inputs MUX are sufficient to implement the proposed MPC. This is a major advantage compared to the approach in [12], whose implementation requires 8 OP-AMPs, two for each inequality.
5.1 Robust Performance Assessment
[…]
To assess robust performances in the presence of parametric uncertainty, we perform $500$ random runs considering random, uniformly distributed, values for each component in the corresponding admissible interval reported in Table 2.
We study the system response to both load and line variations. Figures 6 and 7 show the response of the system when a step disturbance $i_{o}(t)$ is applied, for the case of ceramic and electrolytic capacitors, respectively. The amplitude of the current step is $I_{o,\mathrm{max}}-5/R_{L}$, where $I_{o,\mathrm{max}}$ is the maximum output current bearable by the circuit, such that the total output current jumps to such worst case value.
Figure 6: Load disturbance: ceramic capacitor
Figure 7: Load disturbance: electrolytic capacitor
[…]
Figure 8: Line disturbance: ceramic capacitor
Figure 9: Line disturbance: electrolytic capacitor
[…]
The line disturbance rejection performances achieve a very reduced settling time, which is, on average, comparable to just a few cycles of the PWM modulation frequency. The undershoot/overshoot strongly depends on the instantaneously drawn current step and the ESR value, being significantly larger for the electrolytic capacitor case. As expected, we also observe a small steady-state tracking error due to the presence of line disturbances, resulting from the linearized model’s nonlinearity involving $V_{\mathrm{in}}$. However, this effect is always comparable to, or less than, the voltage ripple.
Finally, we observe that when considering the electrolytic capacitor, the undershoot becomes more pronounced with elevated capacitor ESR; however, the settling time improves. In all simulations, three cycles are sufficient to restore steady-state cyclostationary operation. At the same time, the electrolytic capacitor highlights the effect of model linearization in inducing a steady-state tracking error.
5.2 Circuital Simulations
Figure 10: Circuit schematic simulation at the top-level.
This section presents numerical results from circuit simulations conducted using the LTSpice simulator engine [45]. The simulations are extremely accurate, demonstrating the practical feasibility of the proposed approach in real-world implementations and investigating the effects of non-idealities inherent to the practical implementation of the proposed EMPC. In order to conduct the simulations, we select commercially available devices to implement the system.
The overall circuit schematic is shown in Fig. 10. The power stage of the Buck converter in Fig. 4 is implemented resorting to the synchronous Buck controller LTC7060. The latter drives the high-side and low-side power MOSFETs of the half-bridge, respectively BSC050N04LS and BSC016N04LS, depending on the state of the logic signal provided on the PWM input pin. The Dickson’s charge pump, including $C_{\rm pump}$-$D_{\rm pump}$, serves to drive the gate of the n-channel MOSFET BSC050N04LS, and it takes as input the $V_{\rm cc}=$12\text{\,}\mathrm{V}$\$ input voltage. Both the load current estimator (in Fig. 5) and the generalized adders implementing the unsaturated region control laws (in Fig. 2) exploit the AD8031 OP-AMP. AD8031 is a rail-to-rail OP-AMP that operates with a single-supply voltage and is characterized by an $80\text{\,}\mathrm{MHz}$ gain-bandwidth product. The single-supply comparator used to implement the region selector stage, as shown in Fig. 3, is the LT1721. The inductor current sensing is performed through a $10\text{\,}\mathrm{m\SIUnitSymbolOhm}$ shunt resistance $R_{\rm sense}$, together with the current-sense amplifier AD8410A. The latter offers a 2.2 $2\text{\,}\mathrm{MHz}$ bandwidth (four times higher compared to $f_{\rm sw}$), and it is configured so that its voltage gain is $A_{\rm sense}=20$. Therefore, overall the gain $g_{i_{L}}=R_{\rm sense}A_{\rm sense}=$200\text{\,}\mathrm{m}\mathrm{V}\mathrm{/}\mathrm{A}$\$. The OP-AMPs, the current-sense amplifier, and the comparators are supplied with a $V_{\rm DD}=$5\text{\,}\mathrm{V}$\$ voltage. Finally, the selected 4-channel analog MUX is the ADG5404, operating with the $V_{\rm cc}$ supply voltage.
For each component, the manufacturer’s model has been imported into the LTSpice scheme. This enables accurate transistor-level simulation results, which include the non-idealities of the selected components. Among these, we have the finite gain-bandwidth product, the offset currents and voltages, and slew-rate limitations for the OP-AMPs, delays in the response of comparators, MUX, and logic. The values of all passive components (resistors and capacitors) are rounded to the nearest integer.
We conduct a total of nine simulations, considering three values of supply voltage and three values of resistive load. Figures 11-13 show the load disturbance response in the considered high-fidelity simulation scenario, where the input supply voltage is fixed at the values of $40\text{\,}\mathrm{V}$$50\text{\,}\mathrm{V}$ and $60\text{\,}\mathrm{V}$, respectively. As load disturbance, we consider a $10\text{\,}\mathrm{A}$ current pulse with fixed duration of $0.2\text{\,}\mathrm{ms}$. We compare the response obtained with different loads $R_{L}\in\{$1\text{\,}\mathrm{\SIUnitSymbolOhm}$,$3\text{\,}\mathrm{\SIUnitSymbolOhm}$,$5\text{\,}\mathrm{\SIUnitSymbolOhm}$\$\}. In each condition, the response is compared to that achieved using standard voltage mode control (VMC) design for the considered device, also obtained through accurate LTSpice simulation. For details on the design and implementation of the VMC controller for this case study, we refer the reader to [18].
[…]
Figure 11: Load disturbance response: $V_{\mathrm{in}}=$40\text{\,}\mathrm{V}$\$
Figure 12: Load disturbance response: $V_{\mathrm{in}}=$50\text{\,}\mathrm{V}$\$
Figure 13: Load disturbance response: $V_{\mathrm{in}}=$60\text{\,}\mathrm{V}$\$
[…]
In all runs, we observe a small steady-state tracking error for the proposed EMPC. Such an error, always lower than $50\text{\,}\mathrm{mV}$, arises due to the joint effect of the modeling uncertainty in $R_{L}$, the model linearization error, and the need to approximate the gains defining the EMPC lay $\pi(\cdot)$ rounding resistors values. Conversely, VMC features a very small steady-state tracking error thanks to the inherent integral action in the control loop, which guarantees theoretically null steady-state tracking error. The error that we observe in the VMC case, approximately $10\text{\,}\mathrm{mV}$, is due to the non-ideality of the OP-AMP implementing the corresponding controller. The presence of such an error is the main shortcoming of the proposed approach, and future work will address its reduction or, possibly, elimination.
6 Conclusions
This paper has introduced an analog circuit implementation of explicit Model Predictive Control (MPC) for Buck converters. The proposed solution exploits the piecewise affine structure of explicit MPC to map the control law directly into an analog architecture. Furthermore, state-of-the-art methodologies are adopted to minimize the number of regions, thus enabling real-time operation within fast sampling rates, a reduced number of components for implementation, and avoiding the overhead associated with digital implementations.
We theoretically establish stability guarantees for the closed-loop system using contraction theory to establish a sufficient condition for global exponential stability. The proposed circuit was validated through LTSpice and MATLAB Simulink simulations, which confirmed its ability to deliver robust disturbance rejection and maintain closed-loop performance across a wide range of uncertainties in Buck components.
Future work will investigate strategies to overcome the presence of steady-state tracking error in the presence of modeling uncertainty and the application of the approach to other plants requiring fast sampling, e.g., different DC-DC converter topologies.
Acknowledgment
This manuscript reflects only the authors’ views and opinions; neither the European Union nor the European Commission can be considered responsible for them.